OpenStudy (anonymous):

The number of phone calls that arrive at a phone exchange is often modeled as a poisson random variable. Suppose that on the average there are 24 calls per hour. What is the probability that there is no calls in 5 minutes?

OpenStudy (wasiqss):

c on average there is 60/24 =2.5 calls per min

OpenStudy (anonymous):

how does that help me find the probability there are no calls in 5 minutes?

OpenStudy (wasiqss):

we must evaluate proability of getting a call every min

OpenStudy (wasiqss):

u can do that or i shud do

OpenStudy (anonymous):

you should probably haha

OpenStudy (wasiqss):

y he deleted it :P

OpenStudy (anonymous):

i dont know!

OpenStudy (wasiqss):

now its easy

OpenStudy (wasiqss):

24/60 = number of calls per minute

OpenStudy (wasiqss):

sorry i mis read it

OpenStudy (wasiqss):

0.4 calls per min

OpenStudy (mathmagician):

made a mistake- during 5 minutes there will be (24/60)*5=2 calls, so \[\lambda=2\] and probability that there will be no calls is \[2^{0}*e ^{-2}/0!=0.135\]

OpenStudy (anonymous):

Great, Thanks! Could you answer my other guestion mathmagician?